# Text-centric Alignment for Multi-Modality Learning

### v1

@@ 3.4. Modality Summarization
Compared with text-style translation, which maps the training distribution directly to the inference distribution, modality summarization involves utilizing an LLM to summarize the translated data. This step aims to achieve two main objectives: (1) It facilitates the alignment of different modalities in a closer semantic space and further eliminates the heterogeneity of text representation of diverse modalities. As discussed in the previous section, varying text transformation methods cause gaps in semantic space. TAMML further generalizes the text representation across all modalities with similar linguistic styles and formats. (2) It identifies the connections and interactions among elements of different modalities. This process enhances information quality by promoting interaction, such as generating novel information, highlighting shared core information, and eliminating redundant information.
The modality summarization is also generated by LLM. Our experiment is divided into two steps. First, we extract a sample in $D_{E}$ with the inference modality combination. We pre-define the linguistic styles in the prompt, instructing the LLM to integrate information across diverse modalities to produce a summarized output. Second, the output is merged back to our original prompt to form a demonstration, and then one-shot in-context learning is applied to each sample x in $T(D_{T})$. Each sample may follow the textual structure of the demonstration established in the first step, aiming to reduce heterogeneity. Appendix A Fig. 6 demonstrates the details of the modality summarization prompt.
@@ 3.5. LLM Reasoning Augmentation
First, the LLM was assigned a specific prediction task accompanied by straightforward instructions and examples to leverage its external knowledge while analyzing the concepts and information present in the original textual inputs. Throughout the task, the LLM will make predictions and provide comprehensive explanations for each input sample. By utilizing the outputs derived from this predictive reasoning process, we enhance the original input information and achieve data augmentation. Next, for each reasoning process, the LLM was furnished with pertinent background information and examples related to its prediction tasks. It is subsequently assigned an assistant role to provide informative reasoning and accurately predict either objective values in regression tasks or labels in classification tasks. Appendix A Fig. 7 demonstrates the details of the reasoning augmentation prompt.
@@ 3.6. Downstream Training
This section elucidates the entire flow of training and inference within downstream tasks. The process is shown in Fig. 3. Initially, the raw input contains different modalities, which are each very heterogeneous in expression. First, we transform each modality into text and conduct the text-style translation step. In parallel, the text representation is then subjected to two distinct yet complementary processes: modality summarization and reasoning augmentation. The summarization aligns textual representation in diverse modalities, while the reasoning augmentation step integrates these points with external knowledge, thereby enhancing the overall informational context.
The transformed text representation is then fed into a transformer. More specifically, we adopted Longformer(Beltagy et al., 2020) as the transformer architecture. After mean pooling is applied to aggregate the contextual representations, a Multilayer Perceptron (MLP) is employed to process the pooled features and output the predictions. The loss function employed is contingent on the downstream task.
@@ 4. Experiments
Here, we articulate our hypotheses and address the research questions to evaluate the effectiveness of TAMML . Q1: Under modality mismatch scenarios, is TAMML better than the embedding-based SOTA solutions? Q2 (follow Q1): Is TAMML still effective for situations in which the testing modality has been involved during training? (i.e. training: all modalities, testing: some of the modalities) Q3: Is text representation generally more robust than embedding representation for cross-modality translation? Detailed descriptions of our experimental setup, including model checkpoints, hyperparameters, and dataset specifics, are provided in Appendix B. Additional inquiries, Q4 and Q5, are addressed in Appendix D.
Setup In this section, we primarily shown results from GPT-3.5-turbo as our large language models (LLMs) and GPT-4-vision for image captioning, unless specified otherwise. For additional results involving other LLMs and image caption models, please see Appendix C.2.
@@ 4.1.1. Dataset
PetFinder.my Adoption Prediction (Addison Howard, 2018) examines the predictive factors of pet adoptability, with a particular focus on the rapidity of adoption post-listing. The dataset is a composite of the following modalities:
  • Text modality contains the human-written description of the status of the pet
  • Image modality contains a profile photo for the pet
  • Tabular modality contains basic information, such as gender and age, of the pet.
@@
Airbnb Pricing Prediction (ins, 2023) is a composite of the following modalities that we used for making a regression prediction of housing prices in the Hawaii market:
@@
  • Text modality contains the human-written description of the homestay, the neighborhood description, and the host’s profile.
  • Image modality contains images of the homestay
@@
  • Tabular modality delivers essential details such as location, rating score, and review counts.
Training Testing PetFinder  —  Accuracy $\uparrow$ Airbnb  —  MSE $\downarrow$
SDEdit DDRM Idinvert TAMML SDEdit DDRM Idinvert TAMML
text+image tabular 0.282 0.291 0.279 0.356 0.935 0.600 0.799 0.244
text+tabular image 0.289 0.277 0.286 0.355 0.656 0.778 0.643 0.394
image+tabular text 0.281 0.297 0.279 0.330 0.514 0.565 0.781 0.285
text image+tabular 0.291 0.283 0.289 0.343 1.548 0.914 0.915 0.216
text image 0.289 0.276 0.287 0.350 1.513 0.895 1.010 0.404
text tabular 0.293 0.259 0.277 0.342 1.061 0.824 0.931 0.233
image text+tabular 0.290 0.297 0.284 0.328 0.556 0.530 0.602 0.391
image text 0.288 0.282 0.280 0.324 0.678 0.589 0.759 0.423
image tabular 0.291 0.287 0.284 0.364 0.592 0.538 0.516 0.428
tabular text+image 0.290 0.271 0.285 0.355 0.637 0.675 0.662 0.290
tabular text 0.289 0.265 0.280 0.326 0.569 0.693 0.707 0.309
tabular image 0.289 0.263 0.277 0.357 0.609 0.715 0.615 0.408
Average 0.289 0.279 0.282 0.344 0.822 0.693 0.745 0.335
@@
Table 1. This table presents a detailed comparison, highlighting TAMML ’s performance against all baseline models under modality mismatch scenarios. PetFinder dataset uses accuracy as the key evaluation metric. Airbnb dataset uses Mean Squared Error (MSE) as the key evaluation metric. Avito dataset uses Root Mean Squared Error (RMSE) as the key evaluation metric.
@@ 4.1.2. Competitors
Here, we adopt the embedding-based cross-modality translation solutions. A high-level illustration of conventional embedding-based downstream training is demonstrated in Figure 2. In our experiments, we applied various foundation models, each for different modalities, as an encoder that extracts embedding representations, after which alignment layers are used for learning and concatenating each representation. The details of each foundation model and its usage are in Appendix B.4. Following these modality encoders, we concatenate the representative outputs from all modalities and feed them into a transformer encoder. After fine-tuning, the model generates cross-modality alignment outputs.
In the experiments, we compared TAMML to several zero-shot cross-modality data translation methods built on top of embedding-based downstream model training. We include two perturbation-diffusion-based methods, SDEdit (Meng et al., 2022), DDRM  (Kawar et al., 2022), and one GAN inversion method, In-domain inversion GAN (Idinvert) (Zhu et al., 2020). All these methods include training a generative model (diffusion or a GAN) that can translate the input embedding distribution into a distribution similar to training modalities. At test time, we use the trained generative models to generate synthesis modality embedding (similar to the distribution of training modalities) for various unseen modalities. For diffusion models, we train a score-based generative model (Ho et al., 2020) and a backbone model for every modality combination. In our implementation, we leverage DDIM (Song et al., 2020a) as the pre-trained model for both SDEdit and DDRM. For GAN, we leverage the StyleGAN (Karras et al., 2019) implementation as the backbone model.
@@ 4.1.3. Results
The key findings outlined in Table 1 underscore the superior performance of TAMML , which achieves substantial gains over competing baselines across various modality combinations. These results suggest that the proposed strategy, which integrates LLMs’ in-context learning with foundation models, holds a decisive edge over all existing methods. Specifically, on the PetFinder dataset, TAMML enhances accuracy by an average of approximately 21%, significantly outperforming other methods. Similarly, in the Airbnb dataset, TAMML achieves an average reduction in mean square error of around 54%, dwarfing the maximum 16% error reduction seen with alternative baselines.
@@ 4.2. Q2: Is the Proposed Solution Still Effective when there is no modality mismatching?
This section presents experiments designed to assess the performance of TAMML given all modalities are available during training. This setup evaluates whether TAMML can surpass embedding-based solutions even in scenarios without modality mismatches. The analysis primarily focuses on comparing embedding-based methods that leverage pre-trained foundation models for modality processing.
Besides the two datasets introduced in Section 4.1.1, we have included one additional dataset. Avito Demand Prediction  [^1] https://www.kaggle.com/competitions/avito-demand-prediction/data predicts the likelihood of an ad selling something based on user item and context features. The dataset is a composite of the following modalities:
  • Text modality contains the ad title and description.
  • Image modality contains a profile photo of the item.
  • Tabular modality contains basic information, such as region, city, item category, ad price, etc.
4.2.1. Baseline
In this section, ”embedding-based solution” refers to the process of extracting each input modality as embeddings, rather than text, for the downstream discriminative model from pre-trained foundation models, given that all modalities are known and available at training time. The foundation models utilized to extract these embedding representations are the same as those discussed in Section 4.1.2. During the experiments, these embeddings are concatenated for downstream training.
@@ 4.2.1. Baseline
Emb TAMML Emb TAMML Emb TAMML
tabular 0.297 0.370 0.336 0.208 0.107 0.045
image 0.286 0.360 0.501 0.359 0.097 0.045
text 0.284 0.338 0.764 0.264 0.110 0.046
image+tabular 0.336 0.404 0.298 0.226 0.093 0.044
text+tabular 0.384 0.374 0.224 0.212 0.102 0.043
text+image 0.296 0.383 0.487 0.247 0.096 0.044
all 0.396 0.413 0.236 0.230 0.090 0.043
Average 0.326 0.377 0.407 0.249 0.099 0.044
Table 2. This table presents a detailed comparison, highlighting TAMML ’s performance against embedding-based solutions when the model is trained on all modalities and tested on different subset modalities.
@@ 4.2.2. Results
The key findings outlined in Table 2 underscore the superior performance of TAMML , which still achieves substantial gains over competing baselines across various modality combinations. These results suggest that despite no modality mismatching, our strategy holds a decisive edge over embedding-based methods. Specifically, on the PetFinder dataset, our technique enhances accuracy by an average of approximately 16%, significantly outperforming embedding-based methods. Similarly, in the Airbnb dataset, TAMML achieves a decrease of approximately 38% in mean square error, indicating a significant improvement in prediction accuracy. Moreover, in the Avito dataset, the decrease is even more pronounced, with a reduction of approximately 55% in root mean square error when applying TAMML .
@@ 4.3.2. Results
Table 3. Text representation shows consistently less performance degradation for cross-modality translation when explicitly compared to embedding representation. Both representations are derived from the same Multimodal LLMs for fair comparison. Nevertheless, transforming from image to caption has a slight performance drawdown.
@@ 4.4. Ablation Studies
This section explores the contribution of individual components within TAMML by conducting ablation studies. We incrementally added modules to evaluate their impact on performance, with findings summarized in Table 4.
@@
SDEdit Text +Modality +Text-style +Reasoning
Transformation Summarization Translation Augmentation
text+image tabular 0.282 0.310 0.321 0.361 0.356
text+tabular image 0.289 0.329 0.365 0.374 0.355
image+tabular text 0.281 0.305 0.295 0.331 0.330
text image+tabular 0.291 0.282 0.296 0.345 0.343
text image 0.289 0.293 0.298 0.357 0.350
text tabular 0.293 0.297 0.318 0.311 0.342
image text+tabular 0.290 0.314 0.289 0.330 0.328
image text 0.288 0.306 0.330 0.327 0.324
image tabular 0.291 0.300 0.307 0.365 0.364
tabular text+image 0.290 0.194 0.366 0.351 0.355
tabular text 0.289 0.193 0.306 0.309 0.326
tabular image 0.289 0.196 0.357 0.357 0.357
Average $\pm$ Variance (×$\times 10^{-4}$) $0.289\pm 0.12$ $0.277\pm 25.91$ $0.321\pm 7.84$ $0.343\pm 4.53$ $\textbf{0.344}\pm\textbf{1.97}$
Table 4. Ablation studies on various components of TAMML . Our observations reveal that text transformations significantly enhance performance across all modality combinations except for tabular data, which are in fixed formatted text. The formatting issue is effectively solved by incorporating a summarization module, resulting in a substantial enhancement in performance. Furthermore, the inclusion of both the translation module and the reasoning augmentation module leads to further improvements in overall performance.

### v2

@@ 3.4. Modality Summarization
Compared with text-style translation, which maps the inference distribution directly to the training distribution, modality summarization involves utilizing an LLM to summarize the translated data. This step aims to achieve two main objectives: (1) It facilitates the alignment of different modalities in a closer semantic space. As discussed in the previous section, varying text transformation methods cause gaps in semantic space. TAMML further generalizes the text representation across all modalities with similar linguistic styles and formats. (2) It identifies the connections and interactions among elements of different modalities. This process enhances information quality by promoting interaction, such as generating novel information, highlighting shared core information, and eliminating redundant information.
The modality summarization is also generated by LLMs. Our experiment is divided into two steps. First, we extract a sample in $D_{T}$ with the inference modality combination. We pre-define the linguistic styles in the prompt, instructing the LLM to integrate information across diverse modalities to produce a summarized output. Second, the output is merged back into our original prompt to form a demonstration, and then one-shot in-context learning is applied to each sample x in $T(D_{T})$. Each sample may follow the textual structure of the demonstration established in the first step, aiming to reduce heterogeneity. Figure. 4 demonstrates the details of the modality summarization prompt.
A Real-world Example: Building on the example from Section 3.2, now the input includes two modalities: image and table. We summarize the textual representations from these modalities. Here is how the summarization looks: ”The patient has moderately differentiated adenocarcinoma of the sigmoid colon, causing an obstruction and dilation of the descending colon.
@@ 3.5. LLM Reasoning Augmentation
First, the LLM is assigned a specific prediction task accompanied by straightforward instructions and examples to leverage its external knowledge while analyzing the concepts and information present in the original textual inputs. Throughout the task, the LLM will make predictions and provide comprehensive explanations for each input sample. By utilizing the outputs derived from this predictive reasoning process, we enhance the original input information and achieve data augmentation. Next, for each reasoning process, the LLM is furnished with pertinent background information and examples related to its prediction tasks. It is subsequently assigned an assistant role to provide informative reasoning and accurately predict either objective values in regression tasks or labels in classification tasks. Figure. 4 demonstrates the details of the reasoning augmentation prompt.
A Real-world Example: Building on the example from Section 3.2, now the input includes two modalities: image and table. The current goal is to determine whether a patient requires hospital observation. The results after augmentation are as follows: ”The obstruction in the sigmoid colon can lead to increased risks of bowel perforation, where the colon wall might rupture due to increased pressure. This complication is serious and requires immediate medical intervention.
@@ 3.6. Downstream Training
This section elucidates the entire flow of training and inference within downstream tasks. The process is shown in Figure. 3. Initially, the raw input contains different modalities, each of which is very heterogeneous in expression. First, we transform each modality into text. In the inference phase, we conduct the additional text-style translation step. In parallel, the text representation is then subjected to two distinct yet complementary processes: modality summarization and reasoning augmentation. The summarization aligns textual representations in diverse modalities, while the reasoning augmentation step integrates these points with external knowledge, thereby enhancing the overall informational context.
The transformed text representation is then fed into a transformer. More specifically, we adopt Longformer(Beltagy et al., 2020) as the transformer architecture. After mean pooling is applied to aggregate the contextual representations, a Multilayer Perceptron (MLP) is employed to process the pooled features and output the predictions. The loss function employed is contingent on the downstream task. In this paper, we use Cross Entropy Loss for classification tasks and Mean Squared Error for regression tasks.
Figure 4. Examples of prompt templates for each modules
@@ 4. Experiments
Here, we articulate our hypotheses and address the research questions to evaluate the effectiveness of TAMML . Q1: Under modality mismatch scenarios, is TAMML better than the embedding-based SOTA solutions? Q2 (follow Q1): Is TAMML still effective for situations in which the testing modality has been involved during training? (i.e. training: all modalities, testing: some of the modalities) Q3: Is text representation generally more robust than embedding representation for cross-modality translation?
Setup In this section, we primarily present results from GPT-4-Vision for image captioning, unless specified otherwise. For additional results involving other image caption models, please refer to Table 5. Furthermore, our performance enhancement is all calculated using the formula: (measured performance - baseline performance) / baseline performance.
@@ 4.1.1. Dataset
PetFinder.my Adoption Prediction (Addison Howard, 2018) examines what factors predict how quickly a pet is adopted after being listed. The dataset is a composite of the following modalities:
  • Text: contains the description of the status of the pet
  • Image: contains a profile photo of the pet
  • Tabular: contains basic information, such as gender and breed.
@@
Airbnb Pricing Prediction (ins, 2023) is composed of the following modalities used for making a regression prediction of housing prices:
@@
  • Text: contains the human-written description of the homestay, the neighborhood description, and the host’s profile.
  • Image: contains images of the homestay
@@
  • Tabular: delivers essential details such as location, rating score, and review counts.
Avito Demand Prediction (Guz et al., 2018) predicts the likelihood of an ad selling something based on user item and context features:
  • Text: contains the ad title and description.
  • Image: contains a profile photo of the item.
  • Tabular: contains basic information, such as region, city, item category, etc.
Training Testing SDEdit DDRM Idinvert TAMML Methods
LLaMa 3 8B Mistral 7B Mixtral 8x7B Mixtral 8x22B GPT-3.5
PetFinder — Accuracy $\uparrow$
text+image tabular 0.282 0.291 0.279 0.309 0.301 0.317 0.332 0.348
text+tabular image 0.289 0.277 0.286 0.306 0.322 0.335 0.323 0.380
image+tabular text 0.281 0.297 0.279 0.303 0.304 0.320 0.313 0.355
text image+tabular 0.291 0.283 0.289 0.315 0.304 0.330 0.368 0.344
text image 0.289 0.276 0.287 0.355 0.341 0.307 0.360 0.374
text tabular 0.293 0.259 0.277 0.295 0.286 0.325 0.341 0.357
image text+tabular 0.290 0.297 0.284 0.310 0.322 0.346 0.342 0.341
image text 0.288 0.282 0.280 0.323 0.325 0.329 0.330 0.319
image tabular 0.291 0.287 0.284 0.322 0.302 0.341 0.319 0.348
tabular text+image 0.290 0.271 0.285 0.314 0.309 0.327 0.333 0.360
tabular text 0.289 0.265 0.280 0.295 0.294 0.302 0.317 0.364
tabular image 0.289 0.263 0.277 0.294 0.305 0.338 0.311 0.364
Average 0.289 0.279 0.282 0.312 0.310 0.326 0.332 0.355
Airbnb — MSE $\downarrow$
text+image tabular 0.935 0.600 0.799 0.303 0.371 0.326 0.313 0.367
text+tabular image 0.656 0.778 0.643 0.626 0.466 0.451 0.447 0.508
image+tabular text 0.514 0.565 0.781 0.413 0.325 0.312 0.359 0.332
text image+tabular 1.548 0.914 0.915 0.315 0.368 0.323 0.284 0.421
text image 1.513 0.895 1.010 0.537 0.521 0.439 0.404 0.520
text tabular 1.061 0.824 0.931 0.308 0.348 0.345 0.297 0.448
image text+tabular 0.556 0.530 0.602 0.431 0.368 0.382 0.392 0.395
image text 0.678 0.589 0.759 0.439 0.389 0.375 0.421 0.391
image tabular 0.592 0.538 0.516 0.459 0.452 0.487 0.405 0.414
tabular text+image 0.637 0.675 0.662 0.467 0.347 0.310 0.379 0.280
tabular text 0.569 0.693 0.707 0.481 0.341 0.313 0.339 0.301
tabular image 0.609 0.715 0.615 0.627 0.461 0.431 0.535 0.551
Average 0.822 0.693 0.745 0.451 0.396 0.375 0.381 0.411
Avito — MSE $\downarrow$
text+image tabular 0.103 0.113 0.126 0.045 0.045 0.043 0.041 0.044
text+tabular image 0.130 0.133 0.142 0.048 0.048 0.047 0.048 0.046
image+tabular text 0.113 0.125 0.137 0.045 0.045 0.045 0.043 0.046
text image+tabular 0.124 0.123 0.131 0.046 0.047 0.046 0.046 0.045
text image 0.124 0.122 0.129 0.048 0.050 0.047 0.048 0.047
text tabular 0.127 0.124 0.134 0.045 0.046 0.046 0.046 0.044
image text+tabular 0.123 0.126 0.134 0.044 0.044 0.044 0.044 0.044
image text 0.118 0.124 0.129 0.045 0.046 0.047 0.046 0.045
image tabular 0.119 0.126 0.134 0.044 0.044 0.045 0.043 0.044
tabular text+image 0.128 0.139 0.137 0.046 0.045 0.046 0.044 0.046
tabular text 0.124 0.131 0.138 0.046 0.044 0.047 0.044 0.045
tabular image 0.126 0.137 0.140 0.050 0.047 0.048 0.046 0.048
Average 0.122 0.127 0.135 0.046 0.046 0.046 0.045 0.045
@@
Table 1. This table presents a detailed comparison, highlighting TAMML ’s performance against all baseline models under modality mismatch scenarios. The PetFinder dataset uses accuracy as the key evaluation metric. The Airbnb dataset and the Avito dataset both use Mean Squared Error (MSE) as the key evaluation metric.
4.1.2. Large Language Models
To validate our framework and demonstrate its capability to enhance performance, we experiment with different LLM foundation models, including GPT3.5, and open-source models of different sizes in our architecture.
@@ 4.1.3. Competitors
Here, we adopt the embedding-based cross-modality translation solutions. A high-level illustration of conventional embedding-based downstream training is demonstrated in Figure 2. In our experiments, we applied various foundation models, each designated for different modalities, as encoders that extract embedding representations. Afterward, alignment layers are used for learning and concatenating each representation. Following these modality encoders, we concatenate the representative outputs from all modalities and feed them into a transformer encoder. After fine-tuning, the model generates cross-modality alignment outputs.
In the experiments, we compared TAMML to several zero-shot cross-modality data translation methods built on top of embedding-based downstream model training. We include two perturbation-diffusion-based methods, SDEdit (Meng et al., 2022), DDRM  (Kawar et al., 2022), and one GAN inversion method, In-domain inversion GAN (Idinvert) (Zhu et al., 2020). All these methods include training a generative model (diffusion or a GAN) that can translate the input embedding distribution into a distribution similar to training modalities. At test time, we use the trained generative models to generate synthesized modality embeddings (similar to the distribution of training modalities) for various unseen modalities. For diffusion models, we train a score-based generative model (Ho et al., 2020) and a backbone model for each modality combination. In our implementation, we leverage DDIM (Song et al., 2020a) as the pre-trained model for both SDEdit and DDRM. For GANs, we leverage the StyleGAN (Karras et al., 2019) implementation as the backbone model.
@@ 4.1.4. Results
The key findings outlined in Table 1 underscore the superior performance of TAMML , which achieves substantial gains over competing baselines across various modality combinations and different foundation models. These results suggest that the proposed strategy, which integrates LLMs’ in-context learning with foundation models, holds a decisive edge over all existing methods. Specifically, with the best-performing GPT-3.5 on the PetFinder dataset, TAMML enhances accuracy by an average of approximately 21%, significantly outperforming the best-performing baseline methods. Similarly, in the Airbnb dataset, TAMML achieves an average reduction in mean square error of around 54%, dwarfing the maximum 16% error reduction seen with alternative baselines. Further examination of the differences among various foundation models within the TAMML framework underscored the impact of model size on quality. For instance, Mixtral 8x22B improved accuracy by 7% on the PetFinder dataset compared to Mistral 7B. For complex tasks such as summarization and translation, larger models performed better. However, even smaller models showed improvement compared to baselines in mismatch scenarios.
@@ 4.2. Q2: Is the Proposed Solution Still Effective when there is no modality mismatching?
This section presents experiments designed to assess the performance of TAMML given all modalities are available during training. This setup evaluates whether TAMML can surpass embedding-based solutions even in scenarios without modality mismatches. In this section, our TAMML framework employs the Mixtral 8x7B as the foundation LLM and the baselines used for comparison are the same ones used in Section 4.1.
@@
SDEdit DDRM Idinvert TAMML SDEdit DDRM Idinvert TAMML SDEdit DDRM Idinvert TAMML
tabular 0.282 0.269 0.252 0.338 0.428 0.621 0.732 0.270 0.108 0.123 0.133 0.041
image 0.285 0.286 0.267 0.356 0.566 0.649 0.711 0.486 0.114 0.123 0.136 0.044
text 0.284 0.284 0.274 0.349 0.502 0.601 0.695 0.253 0.113 0.123 0.131 0.044
image+tabular 0.307 0.276 0.256 0.382 0.394 0.556 0.683 0.251 0.118 0.124 0.129 0.042
text+tabular 0.315 0.306 0.283 0.377 0.353 0.470 0.544 0.185 0.124 0.124 0.134 0.041
text+image 0.292 0.286 0.244 0.378 0.489 0.537 0.673 0.212 0.110 0.115 0.125 0.043
all 0.334 0.304 0.281 0.395 0.345 0.463 0.542 0.178 0.109 0.114 0.123 0.042
Average 0.300 0.287 0.265 0.368 0.440 0.557 0.654 0.262 0.112 0.121 0.130 0.042
Table 2. This table presents a detailed comparison, highlighting TAMML ’s performance against embedding-based translation baselines when the model is trained on all modalities and tested on different subset modalities.
@@ 4.2.1. Results
The key findings outlined in Table 2 underscore the superior performance of TAMML , which still achieves substantial gains over competing baselines across various modality combinations. These results suggest that despite no modality mismatching, our strategy holds a decisive edge over embedding-based methods. Specifically, on the PetFinder dataset, our technique enhances accuracy by an average of approximately 22.6%, significantly outperforming the best-performing embedding-based methods. Similarly, in the Airbnb dataset, TAMML achieves a decrease of approximately 40.5% in mean squared error, indicating a significant improvement in prediction accuracy. Moreover, in the Avito dataset, the decrease is even more pronounced, with a reduction of approximately 62.5% in mean squared error when applying TAMML .
@@ 4.3.2. Results
Table 3. Text representation shows consistently less performance degradation for cross-modality translation when explicitly compared to embedding representation. Both representations are derived from the same Multimodal LLMs for fair comparison. Nevertheless, transforming from image to caption has a slight performance reduction.
@@ 4.4. Ablation Studies
This section explores the contribution of individual components within TAMML by conducting ablation studies. We incrementally add modules to evaluate their impact on performance, with findings summarized in Table 4. In this section, our TAMML framework employs the GPT-3.5 as the foundation LLM.
@@
SDEdit Text +Modality +Reasoning +Text-style
Transformation Summarization Augmentation Translation
text+image tabular 0.282 0.310 0.321 0.338 0.348
text+tabular image 0.289 0.329 0.365 0.363 0.380
image+tabular text 0.281 0.305 0.295 0.321 0.355
text image+tabular 0.291 0.282 0.296 0.343 0.344
text image 0.289 0.293 0.298 0.341 0.374
text tabular 0.293 0.297 0.318 0.315 0.357
image text+tabular 0.290 0.314 0.289 0.325 0.341
image text 0.288 0.306 0.330 0.336 0.319
image tabular 0.291 0.300 0.307 0.303 0.348
tabular text+image 0.290 0.194 0.366 0.341 0.360
tabular text 0.289 0.193 0.306 0.327 0.364
tabular image 0.289 0.196 0.357 0.353 0.364
Average $\pm$ Variance (×$\times 10^{-4}$) $0.289\pm 0.12$ $0.277\pm 25.91$ $0.321\pm 7.2$ $0.334\pm 2.5$ $\textbf{0.355}\pm\textbf{2.4}$
Table 4. Ablation studies on various components of TAMML . Our observations reveal that text transformations significantly enhance performance across all modality combinations except for tabular data, which is in fixed formatted text. The formatting issue is effectively solved by incorporating a summarization module, resulting in a substantial enhancement in performance. Furthermore, the inclusion of both the translation module and the reasoning augmentation module leads to further improvements in overall performance.